Suppose you are five times older than your cousin. In two years, you will be three times older than him. Four years after that you will be twice as old. What are the current ages of you and your cousin?
Answer by mananth(16946)

cousinn's age =x
my age = y
y=5x
y-5x=0
II condition
y+2=3(x+2)
y+2=3x+6
y-3x=4
III condition
y+6 = 2(x+6)
y+6=2x+12
y-2x=6
Solve for x & y
x=2, y=10